What is the center and the radius of the equation of the circle below?
x^2+y^2+10y-56=0

Answer by CharlesG2(834)

What is the center and the radius of the equation of the circle below?
x^2+y^2+10y-56=0
x^2+y^2+10y-56=0
x^2+y^2+10y=56 bring 56 over
x^2+y^2+10y+25=56+25 complete square
x^2+y^2+10y+25=81
x^2+(y+5)^2=81
equation of circle is (x-h)^2+(y-k)^2=r^2
center is (0,-5)
radius is 9